17. The difference of two numbers is 2. The difference of their squares is 44. Find the numbers. (sets)
Answer #1

L x and y. Let two numbers are From @ Since, x-y = 2 & x² - y2 = 44 x² - y2 = 44 (x+y)(2-) = 44 (x+y) : 2 = 44 44 put a-y=2 n